OpenStudy (anonymous):

What is the area of a circle with a radius of 3 meters? Use twenty- two sevenths for pi. seven and one-seventh square meters eighteen and six-sevenths square meters twenty-eight and two-sevenths square meters thirty-seven and five-sevenths square meters

OpenStudy (austinl):

\(\large{\bf A=\pi r^2}\) Can you take it from there?

OpenStudy (anonymous):

no i kinda suck at math

OpenStudy (austinl):

Okay, we have these as our givens. \(\large{\bf r=3m}\) \(\large{\bf \pi=\frac{22}{7}}\) Now you just need to plug those into the equation. If you need more help, just let me know.

OpenStudy (anonymous):

need more

OpenStudy (anonymous):

help

OpenStudy (calculusxy):

r^2*pi 3^2*22/7 9*22/7 Hope you can take it from there. Just multiply.

OpenStudy (anonymous):

i suck at fractions

OpenStudy (austinl):

Okay, you just need to plug those into the equation given earlier :) \(\Large{\bf A=\pi r^2\\ A=\frac{22}{7}\times(3)^2\\ A=\frac{22}{7}\times9\\ A=\frac{22}{7}\times\frac{9}{1}\\ A=\frac{22\times9}{7}}\) What does that come out to?

OpenStudy (calculusxy):

Multiply 22*9. Divide the product by 7.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nope didnt work

OpenStudy (calculusxy):

What is 22*9?

OpenStudy (anonymous):

198

OpenStudy (calculusxy):

Divide it by 7.

OpenStudy (anonymous):

28

OpenStudy (calculusxy):

What was the remainder?

OpenStudy (anonymous):

the wat?

OpenStudy (calculusxy):

Or easily , what option has twenty eight in it?

OpenStudy (anonymous):

C.

OpenStudy (calculusxy):

That's your answer :)

OpenStudy (anonymous):

thanks again

OpenStudy (calculusxy):

No problem :)